# Fast Radio Bursts as a probe of gravity on cosmological scales

### v3

@@ 2.2.2 Dispersion measure statistics
fixed values $N_{\nu}$ number of neutrino species 3 Planck Collaboration et al. (2020b)
@@ 3.1 MCMC likelihood
$\Omega_{\mathrm{b}}$ $0.044^{+0.006}_{-0.010}$ $0.049^{+0.011}_{-0.010}$ $0.0482^{+0.0029}_{-0.0025}$ 3.8 (3.8)
$\Omega_{\mathrm{cdm}}$ $0.240\pm 0.023$ $0.259\pm 0.012$ $0.2637\pm 0.0031$ 3.9 (3.3)
$h$ $0.78\pm 0.11$ $0.819^{+0.140}_{-0.094}$ $0.690^{+0.030}_{-0.044}$ 3.1 (2.9)
$n_{\mathrm{s}}$ $0.98\pm 0.13$ $0.904^{+0.089}_{-0.070}$ $0.954^{+0.034}_{-0.028}$ 2.5 (2.0)
$\sum m_{\nu}/\mathrm{eV}$ $<0.758$ $<0.455$ $<0.104$ 4.5 (3.1)
$\log_{10}\!\left(\frac{T_{\mathrm{AGN}}}{\mathrm{K}}\right)$ $7.817^{+0.069}_{-0.059}$ $7.76^{+0.24}_{-0.19}$ $7.7990^{+0.0087}_{-0.0078}$ 25.9 (16.8)
$\sigma_{8}$ $0.798^{+0.032}_{-0.096}$ $0.8097^{+0.0040}_{-0.0056}$ $0.8116\pm 0.0040$ 1.3 (1.4)
$\hat{\alpha}_{\mathrm{B}}$ $<1.53$ $0.83^{+0.31}_{-0.72}$ $<0.618$ 1.6 (1.7)
$\hat{\alpha}_{\mathrm{M}}$ $<1.64$ $<1.20$ 1.4 (1.5)
$\log_{10}\!\left(\frac{k_{\mathrm{s}}}{\mathrm{Mpc}^{-1}}\right)$ $<-1.46$ $<-1.46$ 1.0 (1.6)
@@ 4 Results
We show the 68% and 95% confidence regions for the three cases, i.e. WL and FRBs alone as well as WL $\times$ FRBs in Fig. 5. Overall, FRBs have considerably less constraining power than WL alone due to the overall lesser SNR, as WL from a stage-4 survey have a cumulative SNR which is roughly 10 times higher than the one of the FRB mock survey assumed here. Consequently, also the constraints on modified gravity are generally weaker. However, FRBs impose much smaller uncertainties on the baryon density $\Omega_{\mathrm{b}}h^{2}$ and the baryonic feedback strength $\log_{10}(\frac{T_{\mathrm{AGN}}}{\mathrm{K}})$. The latter has been extensively discussed in our previous analysis (Reischke et al., 2023), hence we will refrain from detailing this effect here. Nevertheless, it should be noted that we expect a majority of the cross-correlation improvement to originate from degeneracy breaking due to baryonic feedback calibration with FRBs. Indeed, when cross-correlating shear and DM, the confidence regions of almost all parameters shrink significantly. The marginalised posterior mean and the 68% confidence intervals are quantified in Tab. 3. After $\log_{10}(\frac{T_{\mathrm{AGN}}}{\mathrm{K}})$, the parameters describing the energy contents of the Universe ($\sum m_{\nu}$, $\Omega_{\mathrm{b}}$, $\Omega_{\mathrm{b}}$) experience the greatest increase in precision. Looking at the $\Omega_{\mathrm{b}}$-$h$ degeneracy apparent in Fig. 5, a separate calibration of $h$ further tightens the obtained constraints. In appendix A, we show the posterior for combining Euclid-CS×$\times$×FRB with the Planck2018 (Planck Collaboration et al., 2020a) temperature anisotropy auto-correlation. This breaks the degeneracy $\Omega_{\mathrm{b}}$-$h$, leading to much tighter constraints. Walters et al. (2018) similarly forecasted competitive constraining power when combining FRBs with distance ladder measurement on the background level. Since FRBs can be considered a late-type cosmological probe, this is a sensible addition for future analyses. A smaller but still considerable amount of gain is seen for the primordial power spectrum scale dependency $n_{\mathrm{s}}$, possibly caused by the comparatively large signal of FRBs for small $k$, at which $n_{\mathrm{s}}$ still determines the slope.
@@
Consequently, the 68% confidence intervals change slightly as well. Even though FRBs still can not constrain modified gravity alone, the improvement of $\hat{\alpha}_{\mathrm{B}}$ from WL auto-correlation to WL-FRB cross-correlation slightly increases to $42.5\%$ and the one of $\hat{\alpha}_{\mathrm{M}}$ to $31.1\%$. Many of the other parameter improvements are comparable to the weak modified gravity case, except for $n_{\mathrm{s}}$ and $\sigma_{8}$ which respectively decrease and increase marginally in strong modified gravity. Therefore, we expect to gain constraining power on modified gravity when cross-correlating FRBs with WL, albeit less sizeable than for baryon-related parameters. Adding CMB temperature anisotropy information has little effect on the modified gravity parameter contours, indicating only weak correlation between $\hat{\alpha}_{\mathrm{B/M}}$ and $h$, $\Omega_{\mathrm{b}}$, $n_{\mathrm{s}}$ (see appendix A).
@@ 5 Application to FRBs with host
In Fig. 7 we show the FRBs used for this analysis together with the best fit model, see Tab. 5 in the Appendix for a comprehensive list. The error includes the covariance contribution from the LSS, from the host and from the Milky Way. Due to the low number of FRBs, we maximize the likelihood with respect to the $\Lambda$CDM parameters and then scan the profiles of $\hat{\alpha}_{\mathrm{M}}$ and $\hat{\alpha}_{\mathrm{B}}$ respectively. This removes the issue of projection effects and allows to get the most stringent constraints from the current data set. Fig. 8 shows this likelihood scan, showing that the posterior is prior driven (see Tab. 2). Note that the sharp cut-off in $\hat{\alpha}_{\mathrm{B}}$ is caused by the introduced emulator boundaries. There seems to be no information regarding $\hat{\alpha}_{\mathrm{M}}$ with the current data. While there is a slight constraint on $\hat{\alpha}_{\mathrm{B}}$, this is not significantly different from the prior. Generally this behaviour is expected as the parameter dependence is only present in the perturbations and thus in the covariance, degrading the information content compared to the model. However, it goes to show that host-identified FRBs are in principle a versatile tool to test gravity, in particular with many more sources expected in the next years (e.g. with DSA-2000, Hallinan et al., 2019)
@@ 6 Summary
    First, we investigated a weak modified gravity case with fiducial linear Planck mass run rate and braiding proportionality constant set to $\hat{\alpha}_{\mathrm{M}}^{\mathrm{fid}}=\hat{\alpha}_{\mathrm{B}}^{\mathrm{fid}}=0.05$. In this case, the forecast is dominated by the $\hat{\alpha}_{\mathrm{M/B}}$ prior edges that exclude negative values. We additionally detect a preference for low values of the screening scale $k_{\mathrm{s}}$. However, the posterior is generally rather unconstrained, which we believe to be a consequence of the fiducial cosmology at which the data was drawn. Nevertheless, by combining 10000 FRBs with redshift information with a Euclid-CS-like survey, the $\hat{\alpha}_{\mathrm{M}}$ and $\hat{\alpha}_{\mathrm{B}}$ 68%-constraints improve by a factor of 1.4 and 1.6 respectively, in comparison to the constraints from Euclid-CS alone.
@@
    We repeat the same analysis with strong fiducial modified gravity case with $\hat{\alpha}_{\mathrm{M}}^{\mathrm{fid}}=\hat{\alpha}_{\mathrm{B}}^{\mathrm{fid}}=1$. Here the posteriors are no longer prior dominated and $k_{\mathrm{s}}$-bias is no longer detectable. The improvement from WL to WL×$\times$×FRB changes to a factor of 1.7, 1.5 and 1.6 for $\hat{\alpha}_{\mathrm{M}}$ $\hat{\alpha}_{\mathrm{B}}$ and $\log_{10}(\frac{k_{\mathrm{s}}}{\mathrm{Mpc}^{-1}})$ respectively. FRBs alone put little constraints on Horndeski theory, so the majority of signal in the cross-correlation with WL likely originates from degeneracy breaking between other cosmological parameters. It has already been established that FRBs are highly sensitive to the intergalactic baryonic component, which is where a majority of the signal resides in this work as well. Nevertheless, a 50% decrease in the error bars makes an implementation of FRBs into future modified gravity investigations worthwhile.
@@
    An analysis of 21 FRBs with host identification where the dependence of modified gravity is expressed in the covariance of the $\mathrm{DM}-z$ relation of the FRBs. The constraints on modified gravity are prior dominated in this case with a small bit of information added for $\hat{\alpha}_{\mathrm{B}}$, consistent with the other findings.
@@ Appendix A Additional constraints from Planck CMB
Furthermore, the Planck likelihood carries no information about the Horndeski screening scale and the baryonic feedback strength, hence we extend the Fisher matrix for $\log_{10}(\frac{T_{\mathrm{AGN}}}{\mathrm{K}})$ and $\log_{10}(\frac{k_{\mathrm{s}}}{\mathrm{Mpc}^{-1}})$ with small values ($10^{-5}$) along the corresponding diagonal, thus implying no intrinsic constraining power on these parameters. The posterior contours for the weak modified gravity case are shown in Fig. A.1. While there is a substantial increase in contraining power on $\Omega_{\mathrm{b}}$, the 68 % confidence region for $h$, and $n_{s}$, the $\hat{\alpha}_{\mathrm{B}}$ ($\hat{\alpha}_{\mathrm{M}}$) decreases only by a factor of 1.25 (1.17). The effect is even less for the strong modified gravity case, where the error margin decreases by a factor of 1.01 (1.04) for $\hat{\alpha}_{\mathrm{B}}$ ($\hat{\alpha}_{\mathrm{M}}$). Therefore, adding CMB temperature fluctuation information has only diminishing returns regarding modified gravity.
@@ Appendix B Accuracy and details of the cosmopower emulator
for each mode, $k$ or $z$, depending on the function. Finally, by taking a histogram of the maximum and mean over all modes for each set of input parameters, a visual representation of how the neural network performs at worst and on average is created, shown in Fig. B.2. At worst, the emulator deviates $\leq 3\%$ and performs much better most of the time. Therefore, we exclude the emulator as a dominating systematic in this work.
@@ Appendix D FRBs used for real life application
In Tab. 5 we summarise all the FRBs used in Section 5:
@@
20220207C 262.38 0.04304 79.3 -2.7807 1.272 Sherman et al. (2024)
@@
20220506D 396.97 0.30039 89.1 -2.7715 1.2711 Sherman et al. (2024)
20220825A 651.24 0.241397 79.7 -2.7786 1.2668 Sherman et al. (2024)
20220920A 314.99 0.158239 40.3 -2.862 1.2378 Sherman et al. (2024)
20221012A 441.08 0.284669 54.4 -2.8149 1.2309 Sherman et al. (2024)
20220914A 631.29 0.1139 47.55 1.7812 1.28 Sherman et al. (2024)
20220509G 269.53 0.0894 46.55 1.7919 1.226 Sherman et al. (2024)
20190608A 339.5 0.1178 204.75 2.6882 -0.1378 CHIME/FRB Collaboration et al. (2021)
20200430A 380.1 0.1608 112.21 0.8675 0.216 Kumar et al. (2021)
20150517A 560.0 0.19273 37.53 -1.6929 0.5783 Spitler et al. (2016)
20191001A 506.92 0.234 24.87 2.502 -0.9555 Shannon et al. (2019)
20190714A 504.7 0.2365 154.47 0.0695 -0.2273 Bhandari et al. (2019)
20190102C 363.6 0.291 17.67 2.4856 -1.3871 Macquart et al. (2020)
20180924B 362.4 0.3212 34.43 2.55 -0.7138 Bannister et al. (2019)
20200906A 577.8 0.3688 91.27 -2.2079 -0.2458 Bhandari et al. (2022)
20190611B 321.4 0.378 17.76 2.4565 -1.3857 Macquart et al. (2020)
20181112A 589.27 0.4755 25.76 2.5717 -0.9245 Prochaska et al. (2019)
20190523A 760.8 0.66 18.76 0.4724 1.2648 Ravi et al. (2019)
Table 5: FRBs used for the analysis presented in Section 5.

### v4

@@ 2.2.2 Dispersion measure statistics
fixed values $N_{\nu}$ number of equal mass neutrino species 3 Planck Collaboration et al. (2020b)
@@ 3.1 MCMC likelihood
$\Omega_{\mathrm{b}}$ $0.0455^{+0.0056}_{-0.011}$ $0.046^{+0.012}_{-0.010}$ $0.0484^{+0.0028}_{-0.0025}$ 4.2 (4.7)
$\Omega_{\mathrm{cdm}}$ $0.245\pm 0.022$ $0.261^{+0.010}_{-0.012}$ $0.2625\pm 0.0032$ 3.5 (3.2)
$h$ $0.78^{+0.14}_{-0.12}$ $0.838^{+0.15}_{-0.060}$ $0.686^{+0.030}_{-0.043}$ 2.9 (2.9)
$n_{\mathrm{s}}$ $0.97\pm 0.13$ $0.864\pm 0.072$ $0.956^{+0.037}_{-0.030}$ 2.2 (2.2)
$\sum m_{\nu}/\mathrm{eV}$ $<0.817$ $<0.647$ $<0.132$ 5.0 (4.1)
$\log_{10}\!\left(\frac{T_{\mathrm{AGN}}}{\mathrm{K}}\right)$ $7.821^{+0.058}_{-0.065}$ $7.75^{+0.22}_{-0.18}$ $7.7995^{+0.0090}_{-0.0070}$ 24.8 (16.8)
$\sigma_{8}$ $<0.812$ $0.8071^{+0.0050}_{-0.0059}$ $0.8120^{+0.0044}_{-0.0039}$ 1.3 (1.4)
$\hat{\alpha}_{\mathrm{B}}$ $<1.53$ $0.76^{+0.23}_{-0.73}$ $<0.641$ 1.5 (1.7)
$\hat{\alpha}_{\mathrm{M}}$ $1.37^{+0.54}_{-1.3}$ $<1.35$ 1.3 (1.5)
$\log_{10}\!\left(\frac{k_{\mathrm{s}}}{\mathrm{Mpc}^{-1}}\right)$ $<-1.52$ $<-1.60$ 1.2 (1.7)
@@ 4 Results
We show the 68% and 95% confidence regions for the three cases, i.e. WL and FRBs alone as well as WL $\times$ FRBs in Fig. 5. Overall, FRBs have considerably less constraining power than WL alone due to the overall lesser SNR, as WL from a stage-4 survey have a cumulative SNR which is roughly 10 times higher than the one of the FRB mock survey assumed here. Consequently, also the constraints on modified gravity are generally weaker. However, FRBs impose much smaller uncertainties on the baryon density $\Omega_{\mathrm{b}}h^{2}$ and the baryonic feedback strength $\log_{10}(\frac{T_{\mathrm{AGN}}}{\mathrm{K}})$. The latter has been extensively discussed in our previous analysis (Reischke et al., 2023), hence we will refrain from detailing this effect here. Nevertheless, it should be noted that we expect a majority of the cross-correlation improvement to originate from degeneracy breaking due to baryonic feedback calibration with FRBs. Indeed, when cross-correlating shear and DM, the confidence regions of almost all parameters shrink significantly. The marginalised posterior mean and the 68% confidence intervals are quantified in Tab. 3. After $\log_{10}(\frac{T_{\mathrm{AGN}}}{\mathrm{K}})$, the parameters describing the energy contents of the Universe ($\sum m_{\nu}$, $\Omega_{\mathrm{b}}$, $\Omega_{\mathrm{cdm}}$) experience the greatest increase in precision. Looking at the $\Omega_{\mathrm{b}}$-$h$ degeneracy apparent in Fig. 5, a separate calibration of $h$ further tightens the obtained constraints. In appendix A, we show the posterior for combining Euclid-CS×$\times$×FRB with the Planck2018 (Planck Collaboration et al., 2020a) temperature anisotropy auto-correlation. This breaks the degeneracy $\Omega_{\mathrm{b}}$-$h$, leading to much tighter constraints. Walters et al. (2018) similarly forecasted competitive constraining power when combining FRBs with distance ladder measurement on the background level. Since FRBs can be considered a late-type cosmological probe, this is a sensible addition for future analyses. A smaller but still considerable amount of gain is seen for the primordial power spectrum scale dependency $n_{\mathrm{s}}$, possibly caused by the comparatively large signal of FRBs for small $k$, at which $n_{\mathrm{s}}$ still determines the slope.
@@
Consequently, the 68% confidence intervals change slightly as well. Even though FRBs still can not constrain modified gravity alone, the improvement of $\hat{\alpha}_{\mathrm{B}}$ from WL auto-correlation to WL-FRB cross-correlation slightly increases to $42.5\%$ and the one of $\hat{\alpha}_{\mathrm{M}}$ to $34.6\%$. Other parameter improvements are comparable to the weak modified gravity case, with the slight exception of $\sum m_{\nu}$ and $\Omega_{\rm cdm}$ where the improvement decreases in strong modified gravity, and $\Omega_{\rm b}$ where it increases. Therefore, we expect to gain constraining power on modified gravity when cross-correlating FRBs with WL, albeit less sizeable than for baryon-related parameters. Adding CMB temperature anisotropy information has little effect on the modified gravity parameter contours, indicating only weak correlation between $\hat{\alpha}_{\mathrm{B/M}}$ and $h$, $\Omega_{\mathrm{b}}$, $n_{\mathrm{s}}$ (see appendix A).
@@ 5 Application to FRBs with host
In Fig. 7 we show the FRBs used for this analysis together with the best fit model, see Tab. LABEL:tab:app_frb_list in the Appendix for a comprehensive list. The error includes the covariance contribution from the LSS, from the host and from the Milky Way. Due to the low number of FRBs, we maximize the likelihood with respect to the $\Lambda$CDM parameters and then scan the profiles of $\hat{\alpha}_{\mathrm{M}}$ and $\hat{\alpha}_{\mathrm{B}}$ respectively. This removes the issue of projection effects and allows to get the most stringent constraints from the current data set. Fig. 8 shows this likelihood scan, showing that the posterior is prior driven (see Tab. 2). Note that the sharp cut-off in $\hat{\alpha}_{\mathrm{B}}$ is caused by the introduced emulator boundaries. There seems to be no information regarding $\hat{\alpha}_{\mathrm{M}}$ with the current data. While there is a slight constraint on $\hat{\alpha}_{\mathrm{B}}$, this is not significantly different from the prior. Generally this behaviour is expected as the parameter dependence is only present in the perturbations and thus in the covariance, degrading the information content compared to the model. However, it goes to show that host-identified FRBs are in principle a versatile tool to test gravity, in particular with many more sources expected in the next years (e.g. with DSA-2000, Hallinan et al., 2019)
@@ 6 Summary
    First, we investigated a weak modified gravity case with fiducial linear Planck mass run rate and braiding proportionality constant set to $\hat{\alpha}_{\mathrm{M}}^{\mathrm{fid}}=\hat{\alpha}_{\mathrm{B}}^{\mathrm{fid}}=0.05$. In this case, the forecast is dominated by the $\hat{\alpha}_{\mathrm{M/B}}$ prior edges that exclude negative values. We additionally detect a preference for low values of the screening scale $k_{\mathrm{s}}$. However, the posterior is generally rather unconstrained, which we believe to be a consequence of the fiducial cosmology at which the data was drawn. Nevertheless, by combining 10000 FRBs with redshift information with a Euclid-CS-like survey, the $\hat{\alpha}_{\mathrm{M}}$ and $\hat{\alpha}_{\mathrm{B}}$ 68%-constraints improve by a factor of 1.3 and 1.5 respectively, in comparison to the constraints from Euclid-CS alone.
@@
    We repeat the same analysis with strong fiducial modified gravity case with $\hat{\alpha}_{\mathrm{M}}^{\mathrm{fid}}=\hat{\alpha}_{\mathrm{B}}^{\mathrm{fid}}=1$. Here the posteriors are no longer prior dominated and $k_{\mathrm{s}}$-bias is no longer detectable. The improvement from WL to WL×$\times$×FRB changes to a factor of 1.7, 1.5 and 1.7 for $\hat{\alpha}_{\mathrm{M}}$ $\hat{\alpha}_{\mathrm{B}}$ and $\log_{10}(\frac{k_{\mathrm{s}}}{\mathrm{Mpc}^{-1}})$ respectively. FRBs alone put little constraints on Horndeski theory, so the majority of signal in the cross-correlation with WL likely originates from degeneracy breaking between other cosmological parameters. It has already been established that FRBs are highly sensitive to the intergalactic baryonic component, which is where a majority of the signal resides in this work as well. Nevertheless, a 40% decrease in the error bars makes an implementation of FRBs into future modified gravity investigations worthwhile.
@@
    We conduct an analysis of 71 FRBs with host identification where the dependence of modified gravity is expressed in the covariance of the $\mathrm{DM}-z$ relation of the FRBs. The constraints on modified gravity are prior dominated in this case with a small bit of information added for $\hat{\alpha}_{\mathrm{B}}$, consistent with the other findings.
@@ Appendix A Additional constraints from Planck CMB
Furthermore, the Planck likelihood carries no information about the Horndeski screening scale and the baryonic feedback strength, hence we extend the Fisher matrix for $\log_{10}(\frac{T_{\mathrm{AGN}}}{\mathrm{K}})$ and $\log_{10}(\frac{k_{\mathrm{s}}}{\mathrm{Mpc}^{-1}})$ with small values ($10^{-5}$) along the corresponding diagonal, thus implying no intrinsic constraining power on these parameters. The posterior contours for the weak modified gravity case are shown in Fig. A.1. While there is a substantial increase in contraining power on $\Omega_{\mathrm{b}}$, the 68 % confidence region for $h$, and $n_{s}$, the $\hat{\alpha}_{\mathrm{B}}$ ($\hat{\alpha}_{\mathrm{M}}$) decreases only by a factor of 1.23 (1.15). The effect is even less for the strong modified gravity case, where the error margin decreases by a factor of 1.01 (1.04) for $\hat{\alpha}_{\mathrm{B}}$ ($\hat{\alpha}_{\mathrm{M}}$). Therefore, adding CMB temperature fluctuation information has only diminishing returns regarding this modified gravity parametrisation.
@@ Appendix B Accuracy and details of the cosmopower emulator
for each mode, $k$ or $z$, depending on the function. Finally, by taking a histogram of the maximum and mean over all modes for each set of input parameters, a visual representation of how the neural network performs at worst and on average is created, shown in Fig. B.2. At worst, the emulator deviates $\lesssim 3\%$ and performs much better most of the time. Therefore, we exclude the emulator as a dominating systematic in this work.
@@ Appendix D FRBs used for real life application
In Tab. LABEL:tab:app_frb_list we summarise all the FRBs used in Section 5 sorted by ascending redshift:
Table 5: FRBs used for the analysis presented in Section 5.
@@
20230708A 411.51 0.105 50.0 -2.7889 -0.9661 Shannon & Uttarkar (2023b)
20191106C 332.2 0.10775 25.0 -2.9094 0.7505 CHIME/FRB Collaboration et al. (2021)
20220914A 631.28 0.1139 55.2 -2.8134 1.28 Sherman et al. (2024)
20190608 339.5 0.11778 37.2 -2.7529 -0.1378 Zwaniga & Kaspi (2021)
20190110C 221.6 0.12244 37.1 -2.8515 0.7233 Zwaniga & Kaspi (2021)
20240310A 601.8 0.127 36.0 -3.1182 -0.7756 Shannon & Uttarkar (2024b)
20240213A 357.4 0.1185 32.1 -2.9482 1.2929 Law (2024b)
20230628A 345.15 0.1265 30.83 -2.9475 1.2616 Law (2024b)
20210807D 251.3 0.1293 121.2 -2.7934 -0.0133 James et al. (2022)
20240114A 527.7 0.13 38.82 -2.767 0.0756 Abbott (2024)
20210410D 578.78 0.1415 56.2 -2.7622 -1.3844 Caleb et al. (2023)
20231226A 329.9 0.1569 145.0 -2.9608 0.1066 Shannon & Uttarkar (2024a)
20230526A 361.4 0.157 50.0 -3.1157 -0.9201 Shannon & Uttarkar (2023a)
20220920A 314.99 0.158239 40.3 -2.862 1.2378 Sherman et al. (2024)
20200430 380.25 0.16 27.0 -2.8743 0.216 Kumar et al. (2021b)
20210603A 500.147 0.177 40.0 -3.1296 0.3705 Leung & Collaboration (2022)
20220529A 246.0 0.1839 30.92 -3.1194 0.3601 Li et al. (2025)
20220725A 290.4 0.1926 31.0 -2.7305 -0.6281 Shannon (2022b)
20121102 557.0 0.1927 188.0 -3.045 0.5783 Petroff (2020)
20221106A 343.8 0.2044 35.0 -3.0756 -0.4463 Shannon (2022d)
20240215A 549.5 0.21 42.8 -2.8292 1.2258 Law (2024b)
20210117A 728.95 0.214 34.4 -2.746 -0.2819 Kumar et al. (2021a)
20221027A 452.5 0.229 40.59 -2.9893 1.2584 Law (2023)
20191001 507.9 0.234 44.7 -2.7654 -0.9555 Shannon et al. (2019)
20190714 504.13 0.2365 39.0 -2.9275 -0.2273 Bhandari et al. (2019)
20221101B 490.7 0.2395 192.35 -2.7434 1.2336 Law (2023)
20190520B 1210.3 0.241 113.0 -2.8617 -0.197 Niu et al. (2021)
20220825A 651.24 0.241397 79.7 -2.7786 1.2668 Sherman et al. (2024)
20191228 297.5 0.2432 33.0 -2.7408 -0.5165 Petroff (2020)
@@
20221113A 411.4 0.2505 115.37 -3.0585 1.2271 Law (2024b)
20220831A 1146.25 0.262 187.94 -2.7475 1.2258 Law (2023)
20231123B 396.7 0.2625 33.81 -2.8594 1.2354 Law (2024b)
20230307A 608.9 0.271 29.47 -2.9347 1.2513 Law (2024b)
20221116A 640.6 0.2764 196.17 -3.1169 1.2681 Law (2024b)
20220105A 580.0 0.2785 22.0 -2.8986 0.3921 Shannon & Kumar (2022)
20210320C 384.8 0.2797 42.0 -2.9037 -0.2814 Shannon (2023)
20221012A 441.08 0.284669 54.4 -2.8149 1.2309 Sherman et al. (2024)
20240229A 491.15 0.287 29.52 -2.9438 1.2335 Law (2024a)
20190102 364.5 0.2913 57.3 -2.7664 -1.3871 Macquart et al. (2020)
20220506D 396.97 0.30039 89.1 -2.7715 1.2711 Sherman et al. (2024)
20230501A 532.5 0.301 180.18 -2.746 1.2378 Law (2024b)
20180924 361.42 0.3214 40.5 -2.7622 -0.7138 Bannister et al. (2019)
20230626A 451.2 0.327 32.51 -2.8674 1.2415 Law (2024b)
20180301 536.0 0.3304 152.0 -3.0331 0.0815 Petroff (2020)
20231220A 491.2 0.3355 44.55 -2.9974 1.2856 Law (2024b)
20211203C 635.0 0.3439 63.0 -2.9036 -0.5477 Shannon (2023)
20220208A 437.0 0.351 128.79 -2.7663 1.2224 Law (2023)
20220726A 686.55 0.361 111.42 -3.0556 1.2205 Law (2023)
20230902A 440.1 0.3619 34.0 -3.0809 -0.8261 Shannon & Uttarkar (2023c)
20200906 577.8 0.3688 36.0 -3.0881 -0.2458 Bhandari et al. (2022)
20240119A 483.1 0.37 30.98 -2.8804 1.2499 Law (2024b)
20220330D 468.1 0.3714 57.83 -2.9511 1.2279 Law (2023)
20190611 321.4 0.378 43.67 -2.7684 -1.3857 Macquart et al. (2020)
20220501C 449.5 0.381 31.0 -2.7316 -0.5671 Law (2023)
20220204A 612.2 0.4 46.02 -2.8225 1.2169 Law (2023)
20230712A 586.96 0.4525 30.93 -2.9469 1.2664 Law (2024b)
20181112 589.27 0.4755 102.0 -2.7607 -0.9245 Prochaska et al. (2019)
@@
20190523 760.8 0.66 37.0 -2.9007 1.2648 Ravi et al. (2019)
20240123A 1462.0 0.968 113.01 -3.0622 1.2557 Law (2024b)
20221029A 1391.05 0.975 36.4 -2.9764 1.2645 Law (2023)
20220610A 1457.624 1.016 31.0 -2.7331 -0.5849 Shannon (2022a)
20230521B 1342.9 1.354 209.66 -2.7331 1.2416 Law (2024b)